# Bound on entanglement in neural quantum states

### v1

@@ Introduction.
In tensor network-based states such as matrix product states (MPS) [1, 2, 3, 4, 5], this tradeoff is well understood. While they are amenable to very efficient algorithms exploiting their linear structure, their expressive power is constrained by the area law of entanglement. MPS faithfully describe ground states of gapped one-dimensional systems with short-range correlations, but cannot efficiently represent critical states or states with volume law entanglement. Knowing these limitations has been crucial: it underpins our understanding of their success for gapped ground states [6, 7, 8] and points to where new methods are needed [4, 5, 9, 10].
@@
Neural quantum states (NQS), introduced by Carleo and Troyer [11], have emerged as an alternative variational class based on neural networks. They capitalize on the remarkable expressivity of neural networks and efficiency of modern optimizers for even millions of parameters [12, 13, 14]. In practice, NQS have proven competitive across condensed matter, quantum chemistry, and nuclear physics [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], often achieving state-of-the-art accuracy and competitive scaling compared to tensor network and quantum Monte Carlo [30, 31, 32, 33, 12, 34]. Yet, despite this empirical progress, our theoretical understanding of NQS remains incomplete. In contrast to MPS, where the area law constraint clearly delineates expressive power, clear constraints are largely lacking for NQS. Identifying such constraints, if they exist, is crucial for conceptual understanding and for deploying targeted solutions.
@@
Due to their variability and flexibility, establishing general constraints for NQS is a difficult task. Moreover, by the universal approximation theorem [35], sufficiently large NQS can in principle approximate any state. Therefore, it is important to refine the scope of the problem. In this work, we choose to restrict the number of scalar nonlinearities, $k$, meaning the number of elementary nonlinear operations in the computational graph. Specifically, we consider families of NQS defined for arbitrary system sizes $n$, but with $k$ growing strictly slower than $O(n/\log n)$; for clarity, we often state results in the regime $k=O(1)$. Restricting to these NQS has two advantages. First, these states are computationally cheaper, typically reducing the cost of variational Monte Carlo (VMC). Second, as we’ll show, bounding $k$ allows us to derive a sharp constraint on entanglement.
@@
We prove that for these NQS, entanglement entropy of a subregion $A$ satisfies $S_{A}=O(\mu\log n)$, where $\mu\leq k+1$ is the effective dimension of an intermediate feature space. Thus, no family of NQS can achieve volume law entanglement, which requires $S_{A}=O(|A|)$ for $|A|=O(n)$, unless $k$ is at least $O(n/\log n)$. This provides a neural quantum state analog of the area law constraint in MPS. For the case $k=O(1)$, we demonstrate analytically and numerically that the scaling in $n$ is tight. The bound is remarkably agnostic to architectural choices, holding across a wide range of NQS.
@@
We emphasize that though our main result is an upper bound, the bound is also saturated: for a variety of cases with $k\sim O(1)$ we find $\log n$ entanglement scaling. This surpasses the area law entanglement capacity of MPS, underscoring the remarkable expressive power of NQS even with very few nonlinearities.
@@
We consider a pure quantum state of $n$ discrete degrees of freedom $\boldsymbol{s}=(s_{1},\ldots,s_{n})$, which we refer to as spins but could also be fermion occupation numbers. The state is fully specified by the wavefunction amplitudes $\langle\boldsymbol{s}|\Psi\rangle=\Psi(\boldsymbol{s})$. We assume $\Psi(\boldsymbol{s})$ is given by a feed-forward neural network  [^1] We assume a feed-forward neural network to mean the computational graph is acyclic, which may omit certain architectures such as recurrent neural networks, but includes multilayer perceptrons, convolutional neural networks, transformers, autoencoders, and others. and first establish some structural properties of $\Psi(\boldsymbol{s})$.
The computation of $\Psi(\boldsymbol{s})$ can generally [12] be described by a directed acyclic graph $(V,E)$ which maps inputs $s_{1},\ldots,s_{n}$ to the output $\Psi(\boldsymbol{s})$. The values of the nodes $v\in V$ can be recursively defined by $v(\boldsymbol{s})=\sigma_{v}(b_{v}+\sum_{(u,v)\in E}w_{u,v}u(\boldsymbol{s}))$, where $\sigma_{v}$ is a scalar nonlinearity and the sum is over directed edges $(u,v)\in E$ pointing to $v$. In the special case of input nodes, we also allow simply $v(\boldsymbol{s})=b_{v}+\boldsymbol{w}_{v}^{T}\boldsymbol{s}$. By acyclic we mean there are no directed cycles in the graph. We’ll assume the network parameters are real but allow the nonlinearities and network output to be complex.
@@
More generally, when the network $\Psi$ consists of $k$ nonlinearities, $\mathcal{G}$ can be found with $\mu\leq k+1$ affine features. We call this feature reduction and provide a full constructive proof in the Supplemental Material [37]. The basic reasoning is as follows. A directed acyclic graph (DAG) has a node ordering such that for any directed edge $(u,v)$ between nodes $u$ and $v$, $u$ precedes $v$ in the ordering. We may proceed inductively on such an ordering of the DAG of $\Psi$. The first node must be a function of a single affine feature, and each subsequent node is a function of the preceding affine features and possibly one more. Including an optional overall affine feature then gives $\mu\leq k+1$.
Intuitively, this means that a feed-forward network with $k$ nonlinearities cannot couple the $n$ input spins through more than $k+1$ independent affine features (i.e. collective variables). Thus, if $k\sim O(1)$, the wavefunction amplitudes are functions of a small number of affine features rather than all $n$ degrees of freedom. This will be the key to bounding entanglement.
Next, we must make some well-behavedness assumptions about the NQS. While different assumptions can be adopted here, we will choose to require the following technical condition: that $\mathcal{G}(t_{1},\ldots,t_{\mu})$ is analytic in a complex neighborhood of each of its real arguments, and bounded on this neighborhood, in a manner independent of $n$. In particular, we assume the domains of the $t_{i}$ are bounded independently of $n$, which sets the scaling of weights to be $\|\boldsymbol{w}\|_{1}=O(1)$. With this assumption, we may use standard results in approximation theory [38, 37] that there exists a multivariable polynomial of degree at most $d$ in each variable that approximates $\mathcal{G}$ uniformly on its domain $\mathcal{D}\subset\mathbb{R}^{\mu}$ to accuracy $O(e^{-\gamma d})$, where $\gamma>0$ is a constant.
[…]
@@
Here, $\boldsymbol{x}=(x_{1},\ldots,x_{\mu})$, $\mathcal{D}$ is the bounded domain of $\mathcal{G}$, and $\beta,\gamma$ are $O(1)$ constants. We are now in a position to state our main result.
@@
Figure 1: (a) Subregion von Neumann entanglement entropy $S_{A}$ for $n=22$ and (b) bipartite von Neumann entanglement entropy $S_{n/2}$ for the Dicke state, Eq. (10), on a logarithmic $x$-axis, corroborating analytically predicted scaling, respectively, up to finite-size effects. (c,d) Subregion and bipartite entanglement entropy for single-nonlinearity, multilayer perceptron, and transformer neural quantum states (SN-NQS, MLP-NQS, T-NQS) on a logarithmic $x$-axis with random parameters and shading indicating one standard deviation. The MLP-NQS has width $3$ and depth 22. For the T-NQS, we chose patches of size 66, stride 55, 4 attention heads, and 2 layers. Each point was averaged over 20 trial wavefunctions.
@@
where $\boldsymbol{w}=(1,\ldots,1)$ and $\sigma(x)$ is any nonlinearity which evaluates to $\delta_{x0}$ on integers $x$ (here we’ll assume $s_{i}=\pm 1$). Let $A$ be an extensive subregion, $|A|=cn$ with $c\sim O(1)$. Then, for large $n$, $S_{A}=(c/2)\log n+O(1)$ [37]. This shows the asymptotic scaling of our bound is optimal. Moreover, this shows logarithmic entanglement scaling can be achieved by NQS with even a single nonlinearity. We plot finite-size numerical results for this state in Fig. 1.
@@
Next, we show numerically that the scaling of our bound with $n$ appears to be saturated by a wide variety of feed-forward NQS, within accessible sizes. For this purpose, we choose three architectures. First, a single-nonlinearity neural quantum state (SN-NQS), whose log-amplitudes, up to normalization, are given by $\sigma(\boldsymbol{w}^{T}\boldsymbol{s}+b)$ where $\sigma=\sigma_{1}+i\sigma_{2}$ and $\sigma_{1,2}$ are real nonlinearities. Second, a multilayer perceptron neural quantum state (MLP-NQS), which processes inputs $\boldsymbol{s}$ through a vanilla neural network of width $w_{0}$ and depth $d_{0}$ before linear projections to the real and imaginary parts of the log-amplitude. And third, a transformer neural quantum state (T-NQS), which splits $\boldsymbol{s}$ into $O(1)$ patches, processes them using a sparse self-attention mechanism [43, 44], concatenates the resulting vectors, and linearly projects onto the real and imaginary parts of the log-amplitude  [^2] Note the transformer architecture includes vector nonlinearities. These can be decomposed into combinations of scalar nonlinearities and linear operations. The transformer architecture with $O(1)$ patches, heads, and layers will thus still have $O(1)$ scalar nonlinearities and thus belong to our framework.. These are shown in Fig. 1.
For each of these architectures, we calculate the entanglement entropy of a subset $A$ of $n$ qubits. In each case, the total number of nonlinearities $k$ remains fixed and $O(1)$. We treat free parameters as independent random variables drawn from normal distributions and choose $\tanh$ nonlinearities (see [37] for further details). In Fig. 1c, where we fix $n$ and vary $|A|$, we find that entanglement in each ensemble of states grows logarithmically. This saturates our bound, which constrains $S=O(\log n)$ for $O(1)$ nonlinearities. In Fig. 1d, we fix $|A|=n/2$ and vary $n$, finding scaling in $n$ clearly consistent with our bound for SN-NQS and MLP-NQS.
@@
In this work, we have provided an upper bound on entanglement in feed-forward neural quantum states with $k$ elementary scalar nonlinear operations acting on $n$ spins or lattice fermion degrees of freedom, namely $S=O(k\log n)$ as $n$ tends to infinity. Analytical and numerical examples confirm that this upper bound is saturated for many generic types of NQS. While our main contribution is an upper bound, our work also affirms the substantial expressive power of NQS: logarithmic entanglement can be achieved even with very few nonlinearities. In contrast, logarithmic entanglement scaling in MPS requires a bond dimension growing polynomially with system size.
Rigorous characterizations of NQS expressivity have appeared in restricted settings [34, 12, 33, 46, 48], alongside numerous empirical studies of expressivity [49, 50, 31, 51, 52]. Our result complements these works by providing a largely architecture-agnostic upper bound on entanglement capacity, an aspect of expressivity, and capturing how the number of nonlinearities acts as a fundamental bottleneck.
Our work suggest several interesting directions for the future. Relaxing our analysis beyond feed-forward networks to include, for instance, recurrent neural networks will clarify how neural network topology influences entanglement capacity. We expect our analyticity assumptions can be relaxed, e.g. due to similar numerical results on ReLU networks [37]. Finally, our work raises the possibility that NQS with $O(1)$ nonlinearities and linear-in-$n$ computational scaling could efficiently capture quantum critical states or quantum dynamics to times $t\sim\log n$ [37].

### v2

@@ Introduction.
In tensor network-based states such as matrix product states (MPS) [50, 36, 29, 46, 8], this tradeoff is well understood. While they are amenable to very efficient algorithms, their expressive power is constrained by an area law upper bound on entanglement. MPS faithfully describe ground states of gapped one-dimensional systems with short-range interactions, but cannot efficiently represent critical states or states with volume law entanglement. Knowing these limitations has been crucial: it underpins our understanding of their success for gapped ground states [18, 3, 23] and points to where new methods are needed [45, 47].
@@
Neural quantum states (NQS), introduced by Carleo and Troyer [5], have emerged as an alternative variational class based on neural networks. They capitalize on the remarkable expressivity of neural networks and efficiency of modern optimizers for a huge number of parameters [38, 25, 34]. In practice, NQS have proven competitive across condensed matter, quantum chemistry, and nuclear physics [26, 6, 7, 35, 42, 14, 33, 20, 32, 52, 2, 16, 12], often achieving state-of-the-art accuracy and competitive scaling compared to tensor network and quantum Monte Carlo [15, 10, 22, 13, 9]. Yet, despite this empirical progress, our theoretical understanding of NQS remains nascent. In contrast to MPS, where our understanding of expressivity constraints is mature, clear constraints are largely lacking for NQS. Identifying such constraints, if they exist, is crucial for conceptual understanding and for deploying targeted solutions.
@@
Due to their variability and flexibility, establishing general constraints for NQS is a difficult task. Moreover, by the universal approximation theorem [21], sufficiently large NQS can in principle approximate any state. Therefore, it is important to refine the scope of the problem. In this Letter, we choose to restrict the number of scalar nonlinearities $k$ in the computational graph, i.e. the number of elementary nonlinear operations or neurons. Specifically, we consider families of NQS defined for arbitrary system sizes $n$, but with $k$ growing strictly slower than $O(n/\log n)$; for clarity, we often state results in the regime $k=O(1)$. Restricting to these NQS has two advantages. First, these states are computationally cheaper, typically reducing the cost of variational Monte Carlo (VMC). Second, as we’ll show, bounding $k$ allows us to derive a sharp constraint on entanglement.
@@
We prove that for these NQS, entanglement entropy of a subregion $A$ satisfies $S_{A}=O(\mu\log n)$, where $\mu\leq k+1$ is the effective dimension of an intermediate feature space. Thus, no family of NQS can achieve volume law entanglement, which requires $S_{A}=O(|A|)$ for $|A|=O(n)$, unless $k$ is at least $O(n/\log n)$, subject to our assumptions. This provides a neural quantum state analog of the area law constraint in MPS. For the case $k=O(1)$, we demonstrate analytically and numerically that the scaling in $n$ is tight. The bound is quite agnostic to architectural choices, holding across a wide range of NQS.
@@
We emphasize that though our main result is an upper bound, the bound is also saturated quite generally: for a variety of cases with $k\sim O(1)$ we find $\log n$ entanglement scaling. This surpasses the area law entanglement capacity of MPS in 1D, underscoring the remarkable expressive power of NQS even with very few nonlinearities.
@@
We consider a pure quantum state of $n$ discrete degrees of freedom $\boldsymbol{s}=(s_{1},\ldots,s_{n})$, which we refer to as spins but could also be fermion occupation numbers. The state is fully specified by the wavefunction amplitudes $\langle\boldsymbol{s}|\Psi\rangle=\Psi(\boldsymbol{s})$. We assume $\Psi(\boldsymbol{s})$ is given by a feed-forward neural network  [^1] We assume a feed-forward neural network to mean the computational graph is directed and acyclic. Most architecture (e.g. multilayer perceptrons, convolutional neural networks, transformers) are included in this definition. Recurrent neural networks are not included in general; however, when evaluated for a finite horizon $T$ they can be unrolled into a DAG. Our entanglement bound then applies to the unrolled DAG with $k$ replaced by the effective number of scalar nonlinearities $k_{\rm eff}$. and first establish some structural properties of $\Psi(\boldsymbol{s})$.
The computation of $\Psi(\boldsymbol{s})$ can generally [38] be described by a directed acyclic graph $(V,E)$ which maps inputs $s_{1},\ldots,s_{n}$ to the output $\Psi(\boldsymbol{s})$. By acyclic we mean there are no directed cycles in the graph. We illustrate an example of such a graph in Fig. 1. The values of the nodes $v\in V$ are defined recursively by
$v(\boldsymbol{s})=\sigma_{v}\left(b_{v}+\sum_{(u,v)\in E}w_{u,v}u(\boldsymbol{s})+\sum_{i}w_{i,v}s_{i}\right)$ (1)
where $\sigma_{v}$ is a scalar nonlinearity, the first sum is over directed edges $(u,v)\in E$ pointing to $v$, and the second sum is over input spins, thus allowing skip connections. The weights and biases $b_{v},w_{u,v},w_{i,v}$ are network parameters for optimization. We’ll assume the network parameters are real but allow the nonlinearities and network output to be complex.
$\boldsymbol{s}$$v\in V$$\Psi(\boldsymbol{s})$
Figure 1: Example of a directed acyclic graph representation of a feed-forward neural quantum state. Input spins $\boldsymbol{s}=(s_{1},\ldots,s_{n})$ (white) are mapped through internal nodes $v\in V$ (green) to an output $\Psi(\boldsymbol{s})$ (red). Internal nodes represent affine combinations followed by scalar nonlinearities, while the output node is an affine combination. We show how the number of nonlinearities $k=|V|$ acts as a bottleneck for entanglement.
@@
More generally, when the network $\Psi$ consists of $k$ nonlinearities, $\mathcal{G}$ can be found with $\mu\leq k+1$ affine features. We call this feature reduction and provide a full constructive proof in the Supplemental Material [1]. The basic reasoning is as follows. A directed acyclic graph (DAG) has a node ordering such that if a directed edge $(u,v)$ connects $u$ and $v$, $u$ precedes $v$ in the ordering. We may proceed inductively on such an ordering of the DAG of $\Psi$. The first node must be a function of a single affine feature, and each subsequent node is a function of the preceding affine features and possibly one more. Including an optional overall affine feature then gives $\mu\leq k+1$.
Intuitively, this means that a feed-forward network with $k$ nonlinearities cannot couple the $n$ input spins through more than $k+1$ independent affine features (or collective variables). Thus if $k$ is small, the wavefunction amplitudes are functions of a small number of affine features rather than all $n$ degrees of freedom. This will be the key to bounding entanglement.
Next, we must make some well-behavedness assumptions about the NQS. Different choices can be adopted here, but the goal is to make a reasonable assumption prohibiting the NQS from “hiding” a large amount of complexity in a few nonlinearities. We will choose to require the following technical condition: that $\mathcal{G}(t_{1},\ldots,t_{\mu})$ is analytic in a complex neighborhood of each of its real arguments, and bounded on this neighborhood, in a manner independent of $n$. In particular, we assume the domains of the $t_{i}$ are bounded independently of $n$, which sets the scaling of weights to be $\|\boldsymbol{w}\|_{1}=O(1)$. We address relaxations of this assumption in the Supplemental Material [1]. With this assumption, we may use standard results in approximation theory [43, 1] that there exists a multivariable polynomial of degree at most $d$ in each variable that approximates $\mathcal{G}$ uniformly on its domain $\mathcal{D}\subset\mathbb{R}^{\mu}$ to accuracy $O(e^{-\gamma d})$, where $\gamma>0$ is a constant.
In fact, our condition can be phrased as a quite natural condition on the nonlinearities: that they are analytic and bounded on a neighborhood of their domains, in a manner independent of $n$. This prohibits any single nonlinearity from hiding a complexity which grows with $n$. Hence our condition encompasses many standard activations used in practice, including $\tanh$, softplus, and polynomials [28].
@@
Here, $\boldsymbol{x}=(x_{1},\ldots,x_{\mu})$, $\mathcal{D}$ is the domain of $\mathcal{G}$, and $\beta,\gamma$ are $O(1)$ constants. We are now in a position to state our main result.
@@
Theorem (4) as stated does not cover nonanalytic activations such as $\mathrm{ReLU}(x)=\max(0,x)$; nevertheless, our numerical experiments suggest the same scaling behavior, and smooth approximations to ReLU (e.g. GeLU) fall within our assumptions. We also note that a recurrent NQS evaluated for a finite horizon $T$ can be unrolled into a DAG with an effective number of scalar nonlinearities $k_{\rm eff}$. If the unrolled DAG satisfies our analyticity assumptions, then the bound applies with $k$ replaced by $k_{\rm eff}$ [1].
Figure 2: (a) Subregion von Neumann entanglement entropy $S_{A}$ for $n=22$ and (b) bipartite von Neumann entanglement entropy $S_{n/2}$ for the Dicke state, Eq. (11), on a logarithmic $x$-axis, corroborating analytically predicted scaling, respectively, up to finite-size effects. (c,d) Subregion and bipartite entanglement entropy for single-nonlinearity, multilayer perceptron, and transformer neural quantum states (SN-NQS, MLP-NQS, T-NQS) on a logarithmic $x$-axis with random parameters and shading indicating one standard deviation. The MLP-NQS has width $3$ and depth 22. For the T-NQS, we chose patches of size $6$, stride 55, 4 attention heads, and 2 layers. Each point was averaged over 20 trial wavefunctions.
@@
where $\boldsymbol{w}=(1,\ldots,1)$ and $\sigma(x)$ is any nonlinearity which evaluates to $\delta_{x0}$ on integers $x$ (here we’ll assume $s_{i}=\pm 1$). Let $A$ be an extensive subregion, $|A|\sim O(n)$. Then, for large $n$, $S_{A}=\frac{1}{2}\log n+O(1)$ [1]. This shows the asymptotic scaling of our bound is optimal. Moreover, this shows logarithmic entanglement scaling can be achieved by NQS with even a single nonlinearity. We plot finite-size numerical results for this state in Fig. 2.
@@
Next, we show numerically that the scaling of our bound with $n$ appears to be saturated by a wide variety of feed-forward NQS, within accessible sizes. For this purpose, we choose three architectures. First, a single-nonlinearity neural quantum state (SN-NQS), whose log-amplitudes, up to normalization, are given by $\sigma(\boldsymbol{w}^{T}\boldsymbol{s}+b)$ where $\sigma=\sigma_{1}+i\sigma_{2}$ and $\sigma_{1,2}$ are real nonlinearities. Second, a multilayer perceptron neural quantum state (MLP-NQS), which processes inputs $\boldsymbol{s}$ through a vanilla neural network of width $w_{0}$ and depth $d_{0}$ before linear projections to the real and imaginary parts of the log-amplitude. And third, a transformer neural quantum state (T-NQS), which splits $\boldsymbol{s}$ into $O(1)$ patches, processes them using a sparse self-attention mechanism [44, 48], concatenates the resulting vectors, and linearly projects onto the real and imaginary parts of the log-amplitude  [^2] Note the transformer architecture includes vector nonlinearities. These can be decomposed into combinations of scalar nonlinearities and linear operations. The transformer architecture with $O(1)$ patches, heads, and layers will thus still have $O(1)$ scalar nonlinearities and belong to our framework.. These are shown in Fig. 2.
For each of these architectures, we calculate the entanglement entropy of a subset $A$ of $n$ qubits. In each case, the total number of nonlinearities $k$ remains $O(1)$. We treat free parameters as independent random variables drawn from normal distributions and choose $\tanh$ nonlinearities (see [1] for further details). In Fig. 2c, where we fix $n$ and vary $|A|$, we find that entanglement in each ensemble of states grows logarithmically. This saturates our bound, which constrains $S=O(\log n)$ for $O(1)$ nonlinearities. In Fig. 2d, we fix $|A|=n/2$ and vary $n$, finding scaling in $n$ clearly consistent with our bound for SN-NQS and MLP-NQS.
@@
So far, we have not imposed a geometry on the $n$-spin system. Let us now consider spins on a $d$-dimensional lattice of linear size $L$, with $n=L^{d}$. In this case, we may apply our bound to area law entangled states, which satisfy $S=O(L^{d-1})$ for extensive subregions. In particular, area law entangled states require NQS with number of nonlinearities $k$ scaling at least as $O(L^{d-1}/\log L)$ under our assumptions. Therefore in dimensions $d>1$ even area law entangled states require a number of nonlinearities growing with system size for accurate NQS representation. This includes, for instance, gapped ground states of local Hamiltonians, which in one dimension are rigorously known and in higher dimensions are widely believed to obey an entanglement area law.
@@
In this work, we have provided an upper bound on entanglement in feed-forward neural quantum states with $k$ elementary scalar nonlinear operations acting on $n$ spins or lattice fermion degrees of freedom, namely $S=O(k\log n)$ as $n$ tends to infinity. Analytical and numerical examples confirm that this upper bound is saturated for many generic types of NQS. While our main contribution is an upper bound on entanglement, our work also affirms the substantial expressive power of NQS: logarithmic entanglement can generally be achieved even with very few nonlinearities. In contrast, logarithmic entanglement scaling in MPS requires a bond dimension growing polynomially with system size.
Our theorem applies to feed-forward (directed acyclic) NQS subject to analyticity assumptions and in which the individual scalar nonlinearities are independent of $n$. It therefore clarifies when volume-law entanglement is and is not compatible with NQS. In particular, volume-law entanglement is permitted by our theorem through several mechanisms: (i) scaling the number of scalar nonlinearities as $k=\Omega(n/\log n)$, as seen in many constructions where the number of neurons grows with $n$ [27, 9]; (ii) employing recurrent or autoregressive architectures, when unrolled for a large enough number of steps; (iii) allowing explicit $n$-dependence in nonlinearities, though this departs from standard machine learning practice; and (iv) incorporating global non-scalar operations such as Slater determinants, which result in an effective $k\sim\mathrm{poly}(n)$. As a concrete example, Ref. [31] learned the volume-law ground state of the quantum Sherrington-Kirkpatrick model using an MLP-NQS with $k\sim n$ neurons. This is consistent with our upper bound, and from our perspective, the choice $k\sim n=\Omega(n/\log n)$ was necessary.
Rigorous characterizations of NQS expressivity have appeared in restricted settings [9, 38, 13, 40, 27, 51], alongside numerous empirical studies of expressivity [41, 30, 10, 31, 49]. Our result complements these works by providing a largely architecture-agnostic upper bound on entanglement capacity, an aspect of expressivity, and capturing how the number of nonlinearities acts as a fundamental bottleneck.
Our work suggest several interesting directions for the future. For example, we expect our analyticity assumptions can be relaxed due to similar numerical results on ReLU networks [1]. In the future, it will be interesting to find upper bounds on expressivity of NQS in terms of other quantities besides entanglement. Moreover, our work raises the possibility that NQS with only $O(1)$ nonlinearities and correspondingly small computational scaling could efficiently capture quantum critical states or quantum dynamics to times $t\sim\log n$ [1].